OpenStudy (anonymous):

help please with 3 questions i will fan and give a medal please i really need help!!

OpenStudy (anonymous):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3

OpenStudy (***[isuru]***):

the amount that katie will gain at the end of a year is \[1200 \times \frac{ 8.5 }{ 100}\] which is 102 then 102 x no.of years = 612 no.of years = 612/102 = 6 years

OpenStudy (anonymous):

6 years is the answer because u need to divide after u multiply and get 612 y.